# LatentExplainer: Explaining Latent Representations in Deep Generative Models with Multimodal Large Language Models

### v6

@@ Abstract.
Deep generative models like VAEs and diffusion models have advanced various generation tasks by leveraging latent variables to learn data distributions and generate high-quality samples. Despite the field of explainable AI making strides in interpreting machine learning models, understanding latent variables in generative models remains challenging. This paper introduces LatentExplainer, a framework for automatically generating semantically meaningful explanations of latent variables in deep generative models. LatentExplainer tackles three main challenges: inferring the meaning of latent variables, aligning explanations with inductive biases, and handling varying degrees of explainability. Our approach perturbs latent variables, interpreting changes in generated data, and uses multimodal large language models (MLLMs) to produce human-understandable explanations. We evaluate our proposed method on several real-world and synthetic datasets, and the results demonstrate superior performance in generating high-quality explanations for latent variables. The results highlight the effectiveness of incorporating inductive biases and uncertainty quantification, significantly enhancing model interpretability.
Explainable AI, Multimodal Large Language Models, Deep Generative Models, Latent Representations
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; May; 2025isbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
To address the aforementioned challenges, we propose LatentExplainer, a novel and generic framework that automatically generates semantically-meaningful explanations of latent variables in deep generative models. Specifically, to explain these variables and work around the black-box nature of the models (addressing Challenge 1), we propose to perturb each latent variable and explain the resulting changes in the generated data. Specifically, we perturb and decode each manipulated latent variable to produce the corresponding sequence of generated data samples. The trend in the sequence is leveraged to reflect the semantics of the latent variable to be explained. To align explanations with the intrinsic nature of the deep generative models (addressing Challenge 2), we design a generic framework that formulates inductive biases on the Bayesian network of latent variable models into textual prompts. These prompts are understandable to large foundation models and humans. To handle the varying degree of explainability in latent variables (addressing Challenge 3), we propose to measure the confidence of the explanations by estimating their uncertainty. This approach assesses whether the latent variables are interpretable and selects the most consistent explanations, ensuring accurate and meaningful interpretation of the latent variables.
@@ 2. Related Work
Diffusion models, proposed by Ho et al. (Ho et al., 2020), generate data by a diffusion process that gradually adds noise to the data and then learns to reverse this process to recover the original data. These models have achieved high-fidelity image generation, surpassing generative adversarial networks (GANs) in quality and diversity. Latent diffusion models allow the model to operate in a lower-dimensional space, which significantly reduces computational requirements while maintaining the quality of the generated samples (Rombach et al., 2022). Advances have made them applicable to text-to-image synthesis (Nichol and Dhariwal, 2021), and audio generation (Kong et al., 2020)
@@ 3. Preliminaries and Problem Formulation
Latent diffusion models further refine the generation process by iteratively refining noise into structured data (Rombach et al., 2022). These models effectively capture the underlying structure of data in a low-dimensional semantic space. Latent variable manipulation of diffusion models aims at transversing the latent variables $\mathbf{z}$ along the semantic latent direction $\mathbf{z_{i}}$. The perturbed vector $\tilde{\mathbf{z}}=\mathbf{z}+\gamma\left[\mathcal{G}(\mathbf{z}+\mathbf{z_{i}})-\mathcal{G}(\mathbf{z})\right]$, where $\gamma$ is a hyper-parameter controlling the strength and $\mathcal{G}$ is a diffusion decoder (Park et al., 2023). An image sequence can then be generated by $\mathcal{G}(\mathbf{\tilde{z}})$. The perturbations in the semantic latent direction lead to semantic changes in the generated image sequence.
@@ 4.2.1. Generic Framework
1 $p$($z_{i}$$\cdot$ ⋅) pattern of change
2 $p(z_{i}\mid$$z_{i^{\prime}}$), $\forall i\neq i^{\prime}$ other variations
@@ 8 $\neq$ change
Table 1. Lookup table for symbol-to-word mapping.
@@ 4.2.4. From Conditional Bias to Prompts
Prompt: If the pattern of change is associated with the group of the property of interest, this image sequence will change as other variations in $[\text{property}_{k}]$.
@@
Prompt: If the pattern of change is not associated with the group of the property of interest, this image sequence will remain constant despite other variations in $[\text{property}_{k}]$.
@@
4:symbols $\leftarrow$ $\mathcal{F}$)
[…]
@@
10:$\mathcal{P}=\pi_{\theta}(\mathcal{F},\mathcal{H},\text{semantics})$triangleright Generate inductive bias prompt
11:$\tilde{z}\leftarrow\textsc{ManipulateLatent}(z,\mathcal{F},\mathcal{P})$triangleright Latent variable perturbation
12:$D_{i}\leftarrow\mathcal{G}(\tilde{z})$triangleright Generate image sequence
$\mathcal{R}\leftarrow\emptyset$triangleright Initialize explanation set
@@
18:$s_{i}=\frac{1}{n-1}\sum_{i\neq j}\textsc{Sim}(r_{i},r_{j})$
19:$\overline{s}=\frac{1}{n\cdot(n-1)}{\sum}_{i=1}^{n}{\sum}_{j=1,i\neq j}^{n}\textsc{Sim}(r_{i},r_{j})$triangleright Compute certainty score
20:$\hat{r}\leftarrow\arg\max_{r_{i}\in\mathcal{R}}s_{i}$
@@ 4.3. Automatic In-context Prompt Generation
By leveraging these three common inductive biases identified in generative models, we can automatically generate prompts $\mathcal{P}$ that align with $\mathcal{F}$ within these three bias using in-context learning. A more detailed example is at Appendix A.
@@ 4.5. Uncertainty-aware Explanations
Uncertainty-aware methods can be applied to large language model responses (Lin et al., 2023) and image explanations (Zhao et al., 2024). To measure the uncertainty of the responses from GPT-4o, we sampled $n$ times from the GPT-4o to generate the responses =$\mathcal{R}=\{r_{1},r_{2},r_{3},...,r_{n}\}.$ The certainty score of the explanation is the average pairwise cosine similarity of the responses $\mathcal{R}$: $1/C\cdot{\sum}_{i=1}^{n}{\sum}_{j=1,i\neq j}^{n}\text{sim}(r_{i},r_{j})$ where $C=n\cdot(n-1)$. Our final explanation $\hat{r}$ is the response that has the highest pairwise cosine similarity with other responses if the latent variable is interpretable. Otherwise, $\hat{r}$ will be “No clear explanation”. The implementation details are in Section 5.5.
@@ 5. Experiment
The experiments are conducted on a 64-bit machine with 24-core Intel 13th Gen Core i9-13900K @ 5.80GHz, 32GB memory and NVIDIA GeForce RTX 4090. We set GPT-4o at $temperature=1$ and $top\_p=1.$ The code is available at  https://github.com/mengdanzhu/LatentExplainer.
@@ 5.1. Dataset
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 5.3. Generative Models under Inductive Biases
5.4. Latent Variable Perturbation
@@ 5.9. Qualitative Evaluation
In Figure 8, we provide the ”young appearance” prompt to Stable Diffusion under conditional bias, and the explanations of all three top latent directions reflect the meaning of youth. The addition of inductive bias prompts can better identify the relation with the property of interest to capture the actual meaning of latent variables. In comparison, the one without the inductive bias prompts in Figure 8, cannot find clear explanations or simply describe the characteristics in the image sequence, lacking an abstract generalization. More qualitative evaluation results can be found in Section  6.
Moreover, we also evaluate the effect of automatically generating inductive bias prompts from the inductive bias formulas within three aforementioned common bias types and the optional information about the symbol $\mathcal{I}$, as shown in Appendix Table 6, and the results are very promising. These examples demonstrate that our generated prompts not only can capture the nuance in the original inductive bias formulas but also help MLLMs generate more precise explanations.
6. More Qualitative Evaluation Results
We provide more qualitative evaluation results for each inductive bias: (1) the disentanglement bias: DDPM in Figure 9, and $\beta$-TCVAE in Figure 10; (2) the combination bias: CSVAE in Figure 11; (3) the conditional bias: Stable Diffusion in Figure 12 and CSVAE in Figure 13.
@@ 7. Conclusion
8. Limitation
In this work, we use GPT-4o as one of our MLLMs to explain the latent variables. Although GPT-4o has improved vision capabilities across most tasks, we still find it incapable of identifying intricate facial features, positions, and orientations. It is necessary to enhance visual understanding to improve the explanation quality for MLLMs.
Appendix A An example of automatically generating prompts from an inductive bias formula
Example: Given a formula $\mathcal{F}$ which is “$p\left(z_{i}\mid t=t_{1}\right)\neq p\left(z_{i}\mid t=t_{2}\right)$ $,\forall t_{1}\neq t_{2}"$, generate a prompt that reflects the inductive bias.
Solution:
Step 1: Extracting mathematical symbols: Extracting all mathematical symbols from the given formula “$p\left(z_{i}\mid t=t_{1}\right)\neq p\left(z_{i}\mid t=t_{2}\right),$ $\forall t_{1}\neq t_{2}"$. Symbols extracted are $\mathcal{S}=\left[\right.$$p\left(z_{i}\mid\cdot\right)"$, “$t=t_{1},t=t_{2},$ $\forall t_{1}\neq t_{2}",``\neq"]$.
[…]
Step 3: Generating a prompt for the given formula $\mathcal{F}$ : We can generate a prompt for the given formula $\mathcal{F}$ using a pretrained LLM $\pi_{\theta}$ with the following prompt that consists of the few-shot examples $\mathcal{H}$ of the three inductive biases, $\mathcal{F}$, and the gathered $\mathtt{semantics}$:
Formula:
$p\left(z_{i}\mid z_{i^{\prime}}=\alpha\right)=p\left(z_{i}\mid z_{i^{\prime}}=\beta\right),\forall i\neq i^{\prime},\alpha\neq\beta$.
$p\left(z_{i}\mid\cdot\right)"$ is “pattern of change"""", “$p(z_{i}\mid z_{i^{\prime}}),\forall i\neq i^{\prime}"$ is “other variations"""", “= """" is “same"""".
Prompt: These two rows of images show the same pattern of change despite other variations. What is the pattern of change? Write in a sentence. If there is no clear pattern, just write “No clear explanation”.
Formula:
[…]
$p\left(z_{i}\mid\cdot\right)"$ is “pattern of change"""", “$p(z_{i}\mid z_{i^{\prime}}),\forall i\neq i^{\prime}"$ is “other variations"""", “G"""" is “a group"""", “"absent"$\in$"∈ " is “associated with"""", “= """" is “same"""", “"absent"$\neq$"≠ " is “change"""".
Prompt: The pattern of change is associated with a group. The first two rows of images show the same pattern of change despite other variations in another group. The pattern of change in the last two rows of images should change given other variations. What is the pattern of change? Write in a sentence. If there is no clear pattern, just write “No clear explanation”.
Formula:
$p\left(z_{i}\mid p_{k}=\alpha\right)\neq p\left(z_{i}\mid p_{k}=\beta\right),\forall z_{i}\in G_{k},\alpha\neq\beta.$
$p\left(z_{j}\mid p_{k}=\alpha\right)=p\left(z_{j}\mid p_{k}=\beta\right),\forall z_{j}\notin G_{k},\alpha\neq\beta.$
$p\left(z_{i}\mid\cdot\right)"$ is “pattern of change"""", “$p_{k}=\alpha,p_{k}=\beta,\forall\alpha\neq\beta"$ is “other variations in the property of interests"""", “G"""" is “a group"""", “"absent"$\in$"∈ " is “associated with"""", “"absent"$\notin$"∉ " is “not associated with"""", “= """" is “same"""", “"absent"$\neq$"≠ " is “change"""".
[…]
Formula:
$p(z_{1}\mid z_{3}=\alpha)=p(z_{1}\mid z_{3}=\beta),\forall\alpha\neq\beta$.
$p\left(z_{1}\mid\cdot\right)"$ is “pattern of change in the first latent variable"""",“$z_{3}=\alpha,z_{3}=\beta,\forall\alpha\neq\beta"$ is “other variations in the third latent variable"""", “="absent"="= " is “same"""". Please give a prompt in plain language.
Output:
Prompt: The pattern of change in the first latent variable remains the same, regardless of variations in the third latent variable. Describe the pattern of change in the first latent variable. If there is no clear pattern, just write “No clear explanation.”
Appendix B Case Study
More specialized cases of automatically generated inductive bias prompts from the inductive bias formulas within the three common types can be found in Table 6.
Figure 10. Example of generated explanations for $\beta$-TCVAE under the disentanglement bias.
Figure 11. Example of generated explanations for CSVAE under the combination bias.
Figure 12. Example of generated explanations for Stable Diffusion under conditional bias.
Figure 13. Example of generated explanations for CSVAE under conditional bias.
Formula Symbol-to-word Mapping Generated Prompt
$p(z_{1}\mid z_{3}=\alpha)=p(z_{1}\mid z_{3}=\beta),\forall\alpha\neq\beta$. $p\left(z_{1}\mid\cdot\right)"$ is “pattern of change in the first latent variable"""",“$z_{3}=\alpha,z_{3}=\beta,\forall\alpha\neq\beta"$ is “other variations in the third latent variable"""", “=” is “same"""". The pattern of change in the first latent variable remains the same, regardless of variations in the third latent variable. Describe the pattern of change in the first latent variable. If there is no clear pattern, just write “No clear explanation.”
$p\left(z_{1}\mid z_{2}=\alpha\right)=p\left(z_{1}\mid z_{2}=\beta\right),\forall z_{1}\in G,z_{2}\in G^{\prime},G\neq G^{\prime},\alpha\neq\beta.$ $p\left(z_{1}\mid z_{3}=\alpha\right)\neq p\left(z_{1}\mid z_{3}=\beta\right),\newline \forall z_{1},z_{3}\in G,\alpha\neq\beta.$ $p\left(z_{1}\mid\cdot\right)"$ is “pattern of change in the first latent variable"""", “$z_{2}=\alpha,z_{2}=\beta,\forall\alpha\neq\beta"$ is “other variations in the second latent variable"""", “$z_{3}=\alpha,z_{3}=\beta,\forall\alpha\neq\beta"$ is “other variations in the third latent variable"""", “G” is “a group”, “$G^{\prime}$” is “another group”, $\in$” is “associated with"""", $\notin$” is “not associated with"""", “=” is “same"""", $\neq$” is “change"""". The pattern of change in the first latent variable is associated with its group but not with another group. This means that the first latent variable remains the same despite variations in the second latent variable (which belongs to another group), but it changes when the third latent variable (which belongs to the same group) changes. Describe the pattern of change in the first latent variable. If there is no clear pattern, write “No clear explanation.”
$p\left(z_{1}\mid p_{k}=\alpha\right)\neq p\left(z_{1}\mid p_{k}=\beta\right),\forall\alpha\neq\beta.$
$\mathcal{I}:p_{k}=\text{age}.$
$p\left(z_{1}\mid\cdot\right)"$ is “pattern of change in the first latent variable"""", “$p_{k}=\alpha,p_{k}=\beta,\forall\alpha\neq\beta"$ is “other variations in the property of interests"""", “$p_{k}$” is “age”, "absent"$\neq$"≠ " is “change"""". The pattern of change in the first latent variable is associated with age. This means that as the property of interest (age) changes, the first latent variable also changes. Describe the pattern of change in the first latent variable. If there is no clear pattern, write “No clear explanation.”
Table 6. Examples of automatically generated inductive bias prompts

### v7

@@ Abstract.
Deep generative models like VAEs and diffusion models have advanced various generation tasks by leveraging latent variables to learn data distributions and generate high-quality samples. Despite the field of explainable AI making strides in interpreting machine learning models, understanding latent variables in generative models remains challenging. This paper introduces LatentExplainer [^1] The code is available at  https://github.com/mengdanzhu/LatentExplainer., a framework for automatically generating semantically meaningful explanations of latent variables in deep generative models. LatentExplainer tackles three main challenges: inferring the meaning of latent variables, aligning explanations with inductive biases, and handling varying degrees of explainability. Our approach perturbs latent variables, interprets changes in generated data, and uses multimodal large language models (MLLMs) to produce human-understandable explanations. We evaluate our proposed method on several real-world and synthetic datasets, and the results demonstrate superior performance in generating high-quality explanations for latent variables. The results highlight the effectiveness of incorporating inductive biases and uncertainty quantification, significantly enhancing model interpretability.
Explainable AI; Multimodal Large Language Models; Deep Generative Models; Latent Representations
journalyear: 2025copyright: ccconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761318isbn: 979-8-4007-2040-6/2025/11ccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
To address the aforementioned challenges, we propose LatentExplainer, a novel and generic framework that automatically generates semantically-meaningful explanations of latent variables in deep generative models. Specifically, to explain these variables and work around the black-box nature of the models, we propose to perturb each latent variable and explain the resulting changes in the generated data. Specifically, we perturb and decode each manipulated latent variable to produce the corresponding sequence of generated data samples. The trend in the sequence is leveraged to reflect the semantics of the latent variable to be explained. To align explanations with the intrinsic nature of the deep generative models, we design a generic framework that formulates inductive biases on the Bayesian network of latent variable models into textual prompts. These prompts are understandable to large foundation models and humans. To handle the varying degree of explainability in latent variables, we propose to measure the confidence of the explanations by estimating their uncertainty. This approach assesses whether the latent variables are interpretable and selects the most consistent explanations, ensuring accurate and meaningful interpretation of the latent variables.
@@ 2. Related Work
Diffusion models, proposed by Ho et al. (Ho et al., 2020), generate data by a diffusion process that gradually adds noise to the data and then learns to reverse this process to recover the original data. These models have achieved high-fidelity image generation, surpassing generative adversarial networks (GANs) in quality and diversity. Latent diffusion models allow the model to operate in a lower-dimensional space, which significantly reduces computational requirements while maintaining the quality of the generated samples (Rombach et al., 2022). Advances have made them applicable to text-to-image synthesis (Nichol and Dhariwal, 2021), and audio generation (Kong et al., 2020).
@@ 3. Preliminaries and Problem Formulation
[…]
@@ 4.2.1. Generic Framework
Table 1. Lookup table for symbol-to-word mapping.
@@
1 pp($z_{i}$ $\mid\cdot$) pattern of change
[…]
@@ 4.2.4. From Conditional Bias to Prompts
Prompt: If the pattern of change is associated with the group of the property of interest, this image sequence will change as other variations in $\text{property}_{k}]$.
@@
Prompt: If the pattern of change is not associated with the group of the property of interest, this image sequence will remain constant despite other variations in $\text{property}_{k}]$.
@@
4:symbols $\leftarrow$ ExtractSymbols($\mathcal{F}$)
5:semantics $\leftarrow\emptyset$
@@
[…]
@@
[…]
@@ 4.3. Automatic In-context Prompt Generation
By leveraging these three common inductive biases identified in generative models, we can automatically generate prompts $\mathcal{P}$ that align with $\mathcal{F}$ within these three bias using in-context learning.
@@ 4.5. Uncertainty-aware Explanations
Uncertainty-aware methods can be applied to large language model responses (Lin et al., 2023) and image explanations (Zhao et al., 2024). To measure the uncertainty of the responses from GPT-4o, we sampled $n$ times from the GPT-4o to generate the responses $\mathcal{R}=\{r_{1},r_{2},r_{3},...,r_{n}\}.$ The certainty score of the explanation is the average pairwise cosine similarity of the responses $\mathcal{R}$: $1/C\cdot{\sum}_{i=1}^{n}{\sum}_{j=1,i\neq j}^{n}\text{sim}(r_{i},r_{j})$ where $C=n\cdot(n-1)$. Our final explanation $\hat{r}$ is the response that has the highest pairwise cosine similarity with other responses if the latent variable is interpretable. Otherwise, $\hat{r}$ will be “No clear explanation”. The implementation details are in Section 5.5.
@@ 5. Experiment
The experiments are conducted on a 64-bit machine with 24-core Intel 13th Gen Core i9-13900K @ 5.80GHz, 32GB memory and NVIDIA GeForce RTX 4090. We set GPT-4o at $temperature=1$ and top_p=1.top\_p=1.
@@ 5.1. Dataset
$\textbf{B}\uparrow$ $\textbf{R}\uparrow$ $\textbf{S}\uparrow$ $\textbf{BS}\uparrow$ $\textbf{BAS}\uparrow$ $\textbf{B}\uparrow$ $\textbf{R}\uparrow$ $\textbf{S}\uparrow$ $\textbf{BS}\uparrow$ $\textbf{BAS}\uparrow$ $\textbf{B}\uparrow$ $\textbf{R}\uparrow$ $\textbf{S}\uparrow$ $\textbf{BS}\uparrow$ $\textbf{BAS}\uparrow$
@@
$\textbf{B}\uparrow$ $\textbf{R}\uparrow$ $\textbf{S}\uparrow$ $\textbf{BS}\uparrow$ $\textbf{BAS}\uparrow$ $\textbf{B}\uparrow$ $\textbf{R}\uparrow$ $\textbf{S}\uparrow$ $\textbf{BS}\uparrow$ $\textbf{BAS}\uparrow$ $\textbf{B}\uparrow$ $\textbf{R}\uparrow$ $\textbf{S}\uparrow$ $\textbf{BS}\uparrow$ $\textbf{BAS}\uparrow$
@@
$\textbf{B}\uparrow$ $\textbf{R}\uparrow$ $\textbf{S}\uparrow$ $\textbf{BS}\uparrow$ $\textbf{BAS}\uparrow$ $\textbf{B}\uparrow$ $\textbf{R}\uparrow$ $\textbf{S}\uparrow$ $\textbf{BS}\uparrow$ $\textbf{BAS}\uparrow$ $\textbf{B}\uparrow$ $\textbf{R}\uparrow$ $\textbf{S}\uparrow$ $\textbf{BS}\uparrow$ $\textbf{BAS}\uparrow$
@@
$\textbf{B}\uparrow$ $\textbf{R}\uparrow$ $\textbf{S}\uparrow$ $\textbf{BS}\uparrow$ $\textbf{BAS}\uparrow$ $\textbf{B}\uparrow$ $\textbf{R}\uparrow$ $\textbf{S}\uparrow$ $\textbf{BS}\uparrow$ $\textbf{BAS}\uparrow$ $\textbf{B}\uparrow$ $\textbf{R}\uparrow$ $\textbf{S}\uparrow$ $\textbf{BS}\uparrow$ $\textbf{BAS}\uparrow$
@@ 5.3. Generative Models under Inductive Biases
5.4. Latent Variable Perturbation Implementations
@@ 5.9. Qualitative Evaluation
In Figure 8, we provide the ”young appearance” prompt to Stable Diffusion under conditional bias, and the explanations of all three top latent directions reflect the meaning of youth. The addition of inductive bias prompts can better identify the relation with the property of interest to capture the actual meaning of latent variables. In comparison, the one without the inductive bias prompts in Figure 8, cannot find clear explanations or simply describe the characteristics in the image sequence, lacking an abstract generalization. More qualitative evaluation results can be found in Figure 9.
@@ 6. Conclusion
7. GenAI Usage Disclosure
We used generative AI tools to improve the grammar and clarity of the writing.
Acknowledgements.
This work was supported by the National Science Foundation (NSF) Grant No. 2414115, No. 2403312, No. 2007716, No. 2007976, No. 1942594, No. 1907805, NIH R01AG089806, and NIH R01CA297856.